Determine the ordered pair solution for the given linear equations. State the ordered pair solution to two decimal places.

1.23y-0.37x=-1.17 and -0.25x-1.57y=2.35

Answer by Alan3354(69443)

1.23y-0.37x=-1.17 times 100 --> -37x + 123y = -117
-0.25x-1.57y=2.35 times -4 --> x + 6.28y = 9.4
x = -6.28y + 9.4
----
Sub for x in the 1st equation.
-37*(-6.28y + 9.4) + 123y = -117
355.36y - 347.8 = -117
355.36y = 230.8
y =~ 0.64948
==============
Sub for y in any equation to find x.
x = -6.28y + 9.4
x =~ 5.32125